OpenStudy (anonymous):

2a^2-4a+2/3a^2-3 Please show all work, and state any excuded values.

OpenStudy (mayankdevnani):

first, use quadratic formula to solve numerator \[\huge 2a^2-4a+2\] \[\huge 2a^2-2a-2a+2\] \[\huge 2a(a-1)-2(a-1)\] \[\huge (a-1)(2a-2)\]

OpenStudy (anonymous):

\[(2a ^{2}-4a+2)/3a ^{2}-3\]=\[2(a ^{2}-2a+1)/3(a ^{2}-1)=2(a-1)^{2}/3(a-1)(a+1)=2(a-1)/3(a+1)\] now if this is equal to 0..then a=1...and anot equal to -1 as the denominator will then become infinite

OpenStudy (mayankdevnani):

then, solve denominator, \[\huge 3a^2-3\] \[\huge 3(a^2-1)\] \[\huge 3(a+1)(a-1)\] Then, fraction becomes, \[\huge \frac{(a-1)(2a-2)}{3(a+1)(a-1)}\] \[\huge \frac{2a-2}{3(a+1)}=\frac{2a-2}{3a+3}\]

OpenStudy (mayankdevnani):

got it @Pyromancer

OpenStudy (mayankdevnani):

@RaGhavv your answer and my answer is same , but only difference is that when simplify your answer then your answer=my answer. And i think that this is the final answer.

OpenStudy (anonymous):

yeah! but he asked for excluded values so i thought it might be an equation...if not ,then also a is not equal to -1 i just added that...and yes both of our answer are indeed correct my friend..:)

OpenStudy (anonymous):

Isn't there a excluded value?

OpenStudy (anonymous):

@RaGhavv @mayankdevnani

OpenStudy (anonymous):

yes there is...a is not equal to that value which makes the denominator infinite..

OpenStudy (anonymous):

ok, and for the final answer i got 2(a-1)/3(a+1)

OpenStudy (mayankdevnani):

you are correct!! @Pyromancer

OpenStudy (anonymous):

yes and so \[3(a+1)\neq0..\]

OpenStudy (anonymous):

i got: \[a\]

OpenStudy (anonymous):

\[a \neq-1\]

OpenStudy (anonymous):

@raghavv

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

yes a is not equal to -1 and rest it can take anyother value..